# Elimination Without Eliminating: Computing Complements of Real Hypersurfaces Using Pseudo-Witness Sets

### v1

@@ Abstract
Many hypersurfaces in algebraic geometry, such as discriminants, arise as the projection of another variety. The real complement of such a hypersurface partitions its ambient space into open regions. In this paper, we propose a new method for computing these regions. Existing methods for computing regions require the explicit equation of the hypersurface as input. However, computing this equation by elimination can be computationally demanding or even infeasible. Our approach instead derives from univariate interpolation by computing the intersection of the hypersurface with a line. Such an intersection can be done using so-called pseudo-witness sets without computing a defining equation for the hypersurface – we perform elimination without actually eliminating. We implement our approach in a forthcoming Julia package and demonstrate, on several examples, that the resulting algorithm accurately recovers all regions of the real complement of a hypersurface.
@@ 1 Introduction
Many hypersurfaces in algebraic geometry arise as the Zariski closure of the projection of another variety. The real complement of such a hypersurface partitions its real ambient space into open regions. This paper presents a new method for computing these regions without requiring a defining equation for the hypersurface. By computing these regions, we mean representing them with a finite set of sample points from which one can determine the number of regions, do membership tests, and extract topological information such as Euler characteristics, for example.
@@
    evaluate the Hessian ${\rm hess}(h)(x)$ at points $x\in\mathbb{C}^{n}$.
@@
$F(a,b,z)=\begin{pmatrix}f(a,b,z)\\[2.15277pt] f^{\prime}(a,b,z)\end{pmatrix}=\begin{pmatrix}z^{2}+az+b\\[2.15277pt] 2z+b\end{pmatrix}$. (1.2)
@@
The real complement ${\mathbb{R}}^{2}\smallsetminus{\mathcal{H}}$ of the hypersurface consists of the two regions $\{(a,b)\in{\mathbb{R}}^{2}\mid h(a,b)>0\}$ and $\{(a,b)\in{\mathbb{R}}^{2}\mid h(a,b)<0\}$. Our algorithm first uses a pseudo-witness set to determine that the degree of $h$ is $\deg h=2$ and constructs the routing function
@@
Using $c=(13,2)$, our algorithm finds four real critical points, and through gradient flow, it determines that three of them belong to one region, whereas the fourth belongs to another region. The result is shown in Figure 1. During that computation, our algorithm never has direct access to the polynomial $h$.
Figure 1: The quadratic discriminant $\{h(a,b)=0\}$ is the black curve in the picture. There are two regions, one above and one below the black curve. Our algorithm represents these regions by critical points (green) of a routing function (with level sets illustrated with colors between magenta and yellow). There is one critical point in the top region and three critical points in the bottom region. The latter three points are connected by gradient flow (the blue curves). The critical points and the flow trajectories were computed without direct access to the polynomial $h=a^{2}-4b$.
The remainder of this paper is organized as follows. Section 2 reviews background on routing functions, gradient roadmaps, and homotopy continuation for the case when the defining polynomial $h$ of $\mathcal{H}$ is known. Section 3 describes how to extend this approach to the case when $h$ is unknown, through the concept of pseudowitness sets, and ends with a high-level description of our method (Algorithm 4). Section 4 presents illustrative examples. The paper concludes in Section 5.
@@ 2 The case of a known defining polynomial
The following collects background information regarding how to compute the real complements of a hypersurface $\mathcal{H}=V(h)$ with known defining polynomial $h$. In particular, we recall the concept of routing functions (Section 2.1) and gradient roadmaps (Section 2.2), as well as homotopy continuation and monodromy solving (Section 2.3).
@@ 2.1 Routing points and routing functions on ${\mathbb{R}}^{n}$
This section is based on [17, Sections 2 and 3]. Let $h\in\mathbb{R}[x_{1},\ldots,x_{n}]$ be a polynomial defining a hypersurface ${\mathcal{H}}=V(h)$. The complement ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$ consists of connected components, that we call regions, on which the sign of $h$ is constant. To compute these regions, we use the routing function
@@
A routing point $y\in Z(r)$ is called nondegenerate if the Hessian matrix ${\rm hess}(r)(y)$ is invertible. The index of a nondegenerate routing point $y$ is the number of positive eigenvalues of ${\rm hess}(r)(y)$. Non-zero eigenvectors of ${\rm hess}(r)(y)$ to positive eigenvalues are called unstable eigenvector directions.
@@
    The norms of $r$, $\nabla r$, and ${\rm hess}(r)$ are bounded on ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$.
@@ 2.2 Gradient flow and connectivity via routing functions
Input: A routing function $r$ on ${\mathbb{R}}^{n}$ associated with a defining polynomial $h$
Output: A partition $\{C_{1},\dots,C_{s}\}$ of the set of routing points $Z(r)$ of $r$ on ${\mathbb{R}}^{n}$ corresponding to the connected components of ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$, where ${\mathcal{H}}=V(h)$.
1ex(1) Find routing points and their indices. Compute all routing points $Z(r)=\{y_{1},\dots,y_{m}\}\subset{\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$ of $r$ and their corresponding indices $i_{1},\dots,i_{m}$.
(2) Initialize adjacency matrix. Set $A=I_{m}$, the $m\times m$ identity matrix.
(3) Populate adjacency matrix via unstable directions.
@@
 foreach unstable eigenvector direction $v$ for ${\rm hess}(r)(y_{j})$ do
      Compute limit routing point from $y_{j}$ in the direction $v$ with respect to $r$, say $y_{w_{+}}$.
      Set $A_{jw_{+}}=A_{w_{+}j}=1$.
      Compute limit routing point from $y_{j}$ in the direction v-v with respect to $r$, say $y_{w_{-}}$.
      Set $A_{jw_{-}}=A_{w_{-}j}=1$.
   end foreach
end for
1ex(4) Transitive closure. Set $M$ to be the transitive closure of $A$.
Partition $Z(r)$ based on the connected components of $M$, say $C_{1},\dots,C_{s}$.
@@ 2.3 Homotopy continuation and monodromy solving
Let $F(a,b,z)$ be as in Example 1.1. We wish to intersect the curve $V(F)\subseteq{\mathbb{C}}^{3}$ with the plane V(a1)V(a-1) by solving $F(1,x_{1},x_{2})=0$. For this, we can consider the homotopy
$H(x,t)=(1-t)F(1,x_{1},x_{2})+\gamma\,t\begin{pmatrix}x_{1}^{2}-1\\[2.0pt] x_{2}-1\end{pmatrix}=0,$
@@
The start points $(1,1)$ and (1,1)(-1,1) are nonsingular solutions of $H(x,1)=0$. Both corresponding paths are trackable and converge respectively to (2,1)(2,-1) and (2,1)(-2,1) of $H(x,0)=0$. This shows that the intersection of the zero set of $F$ with the plane $a=1$ is $\{(2,1,-1),(-2,1,1)\}$.
@@
Recall that in Example 2.4, we computed the intersection of the curve $V(F)$ with the plane V(a1)V(a-1) by solving $F(1,x_{1},x_{2})=0$. We can vary that plane in a monodromy loop: For a loop $\Gamma\colon[0,1]\to{\mathbb{C}}$ with $\Gamma(0)=\Gamma(1)$, we consider now the new homotopy $H(x,t)=F(x_{1},\Gamma(t),x_{2})=0$ with start point (2,1)(2,-1). If $\Gamma(t)=3/4+(1/4)e^{2\pi i(1-t)}$ the corresponding path is trackable and returns to (2,1)(2,-1). In contrast, for the loop $\Gamma(t)=1/4+(3/4)e^{2\pi i(1-t)}$ the path ends at (2,1)(-2,1). Hence, the monodromy action exchanges these two solutions. The reason why the second loop gives a new solution, while first does not, is that the second loop encircles a so-called branch point.
@@ 2.4 Summary of the algorithm for known defining polynomial
(3) Compute regions. Run steps (2)–(5) in Algorithm 1.
@@ 3 The case of an unknown defining polynomial
where $\pi\colon{\mathbb{C}}^{k}\times{\mathbb{C}}^{n-k}\to{\mathbb{C}}^{k}$ is the projection $\pi(p,z)=p$, and the line denotes Zariski closure.
In what follows, we will let $h$ be a polynomial of minimal total degree such that $\mathcal{H}=V(h)$ (this defines $h$ uniquely up to scaling).
The key observation is that in order to carry out step (1)–(3) of Algorithm 2, we do not need to have an explicit expression for $h$. Instead, we just need the following:
@@
  • We need to be able to evaluate $\nabla h$ and ${\rm hess}(h)$ in order to do homotopy continuation in step (2) and gradient flow in step (3).
It turns out that both these pieces of information about $h$ are available through a pseudo-witness set of $\mathcal{H}$. We explain this concept and how it gives the degree in Section 3.1. We then proceed to derive algorithms for evaluating $\nabla r$ and ${\rm hess}(r)$ through the data of a pseudo-witness in Section 3.2, which is the main theoretical contribution of the paper. Building on this, we give a strategy (Algorithm 4) for computing ${\mathbb{R}}^{k}\smallsetminus\mathcal{H}$ in Section 3.3.
@@
It is often useful to consider the complement of the union of $\mathcal{H}$} with another hypersurface $V(g)$ where $g$ is a known polynomial, e.g., when imposing positivity conditions [17, 16]. This situation can be easily handled by simply adding $\log(g)$ to $\log(r)$ above and adjusting $e$ to ensure that $2e>\deg(h)+\deg(g)$.
@@ 3.1 Witness and pseudo-witness sets
It is important to note that the defining equations for $X$ are not used in Definition 3.2; instead, $X$ is represented using $F$. This can potentially introduce multiplicities, which is why we have the second item in the definition. The second item can be rephrased as follows: A witness set for an irreducible component $X$ of $V(F)$ is reduced when the ideal in the primary decomposition of $\langle F\rangle\vcentcolon=\langle f_{1},\ldots,f_{r}\rangle$ corresponding to $X$ is radical. In a reduced witness set, the points in $X\cap\mathcal{M}$ are nonsingular solutions to a square polynomial system in such a way that homotopy continuation can be used to trace the witness set to another witness set $X\cap\mathcal{M}^{\prime}$ for a different slice $\mathcal{M}^{\prime}$. This way, reduced witness sets enable several numerical operations in homotopy continuation, including sampling points on a variety, testing membership, and computing intersections [4, 31, 33, 48, 49, 50]. If a witness set $(F,\mathcal{M},{\mathcal{W}})$ is not reduced, one can utilize deflation to replace the witness system $F$ with a new witness system $F^{\prime}$ so that $(F^{\prime},\mathcal{M},{\mathcal{W}})$ is reduced [32].
@@
We give an example of a non-reduced witness-set. Consider the case $n=2$ and the polynomial $F(x_{1},x_{2})=(x_{1}^{2}+x_{1}^{2}-1)^{2}$. The zero set $X=V(F)$ is the circle. It has dimension $d=1$. The intersection of $V(F)$ with a general line consists of $2$ points.
Take, for instance, the line $\mathcal{M}=V(x_{1})$. Then, $\mathcal{W}=X\cap\mathcal{M}=\{(0,1),(0,-1)\}$. However, $JF(x_{1},x_{2})=4(x_{1}^{2}+x_{1}^{2}-1)\,(x_{1},x_{2})$, so that $JF(0,1)=(0,0)$ and JF(0,1)=(0,0)JF(0,-1)=(0,0) both have rank 0, and not nd=21=1n-d=2-1=1. The witness set $(F,\mathcal{M},\mathcal{W})$ is therefore not reduced. This is caused by the fact that the polynomial $x_{1}^{2}+x_{1}^{2}-1$ defining the circle appears with multiplicity 2 in $F$.
@@
Projecting the coordinates of points in $\mathcal{W}\subset{\mathbb{C}}^{n}$ to the first $k$ coordinates yields $\pi(\mathcal{W})={\mathcal{H}}\cap\mathcal{L}$. In other words, we can compute the intersection of ${\mathcal{H}}\subset{\mathbb{C}}^{k}$ with a general line $\mathcal{L}$ by computing intersection of $X$ with $\mathcal{L}\times\mathcal{L}^{\prime}$. This yields the notion of pseudo-witness set, which we summarize in the following definition.
@@
    Let $X\subset\mathbb{C}^{n}$ be an irreducible variety of dimension $d$, such that ${\mathcal{H}}=\overline{\pi(X)}$ be a hypersurface. A pseudo-witness set for $\mathcal{H}$} is a quadruple $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W})$, where $F$ is a system such that $X$ is an irreducible component of $V(F)$, $\mathcal{L}$ is a general line and $\mathcal{L}^{\prime}$ a general linear space $\mathcal{L}^{\prime}$ of dimension nd1n-d-1, and $\mathcal{W}=X\cap(\mathcal{L}\times\mathcal{L}^{\prime})$.
@@
Since $\pi(\mathcal{W})={\mathcal{H}}\cap\mathcal{L}$, the intersection of the discriminant with $\mathcal{L}$ is given by taking the first two coordinates of these points. The pseudo-witness set $(X,\pi,\mathcal{L}\times{\mathbb{C}},\mathcal{W})$ is reduced.
@@
Figure 3: The left picture illustrates a pseudo-witness set for the red curve $X=V(F)$. While Figure 2 the green linear space $\mathcal{M}$ was general, here it must be a product space: $\mathcal{M}=\mathcal{L}\times\mathbb{C}$, where $\mathcal{L}$ is a general line; see Example 3.6. The pseudo-witness set is then given by the intersection of the red curve $X$ with $\mathcal{L}\times\mathbb{C}$. The projection of $\mathcal{W}$ onto the yellow two-dimensional plane yields $\mathcal{H}$$\cap\mathcal{L}$. The right picture shows the situation within the yellow plane.

### v2

@@ Abstract
Many hypersurfaces in algebraic geometry, such as discriminants, arise as the projection of another variety. The real complement of such a hypersurface partitions its ambient space into open regions. In this paper, we propose a new method for computing these regions. Existing methods for computing regions require the explicit equation of the hypersurface as input. However, computing this equation by elimination can be computationally demanding or even infeasible. Our approach instead derives from univariate interpolation by computing the intersection of the hypersurface with a line. Such an intersection can be done using so-called pseudo-witness sets without computing a defining equation for the hypersurface – we perform elimination without actually eliminating. We implement our approach in a forthcoming Julia package and demonstrate, on several examples, that the resulting algorithm accurately recovers all regions of the real complement of the hypersurface.
@@ 1 Introduction
Many hypersurfaces in algebraic geometry arise as the Zariski closure of the projection of another variety. The real complement of such a hypersurface partitions its real ambient space into open regions. This paper presents a new method for computing these regions without requiring a defining equation for the hypersurface. By computing these regions, we mean representing them with a finite set of sample points from which one can determine the number of regions, perform membership tests, and extract topological information such as Euler characteristics.
@@
    evaluate the Hessian $\operatorname*{Hess}(h)(x)$ at points $x\in\mathbb{C}^{n}$.
@@
$F(a,b,z)=\begin{pmatrix}f(a,b,z)\\[2.15277pt] f^{\prime}(a,b,z)\end{pmatrix}=\begin{pmatrix}z^{2}+az+b\\[2.15277pt] 2z+a\end{pmatrix}$. (1.2)
@@
The real complement ${\mathbb{R}}^{2}\smallsetminus{\mathcal{H}}$ of the hypersurface consists of the two regions $\{(a,b)\in{\mathbb{R}}^{2}\mid h(a,b)>0\}$ and $\{(a,b)\in{\mathbb{R}}^{2}\mid h(a,b)<0\}$. Our algorithm first uses a pseudo-witness set to determine that the degree of $h$ is $2$, picks a random point $c\in{\mathbb{R}}^{2}$, and constructs the routing function
@@
Using $c=(13,2)$, our algorithm finds four real critical points, and through gradient flow, it determines that three of them belong to one region, whereas the fourth belongs to another region; see Figure 1. During this computation, our algorithm never has direct access to the polynomial $h$.
Figure 1: The quadratic discriminant $V(h)$ is the black curve in the picture. There are two regions, one above and one below the black curve. Our algorithm represents these regions by critical points (green) of a routing function (with level sets illustrated with colors between magenta and yellow). There is one critical point in the top region and three critical points in the bottom region. The latter three points are connected by gradient flow (the blue curves). The critical points and the flow trajectories were computed without direct access to the polynomial $h=a^{2}-4b$.
The remainder of this paper is organized as follows. Section 2 reviews background on routing functions, gradient roadmaps, and homotopy continuation for the case when the defining polynomial $h$ of $\mathcal{H}$ is known. Section 3 describes how to extend this approach to the case when $h$ is unknown, through the concept of pseudo-witness sets, and ends with a high-level description of our method (Algorithm 4). In Section 4, we illustrate how our method can be applied to discriminant varieties to explore how the real or positive root count of a parametrized polynomial system varies throughout the parameter space. The paper concludes in Section 5.
@@ 2 The case of a known defining polynomial
This section collects background information regarding how to compute the real complements of a hypersurface $\mathcal{H}=V(h)$ with known defining polynomial $h$. In particular, we recall the concept of routing functions (Section 2.1) and gradient roadmaps (Section 2.2), as well as homotopy continuation and monodromy solving (Section 2.3).
@@ 2.1 Routing points and routing functions on ${\mathbb{R}}^{n}$
This section is based on [17, Sections 2 and 3]. Let $h\in{\mathbb{R}}[x_{1},\ldots,x_{n}]$ be a polynomial defining a hypersurface ${\mathcal{H}}=V(h)$. The complement ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$ consists of connected components, which we call regions, on which the sign of $h$ is constant. To compute these regions, we use the routing function
@@
A routing point $y\in Z(r)$ is called nondegenerate if the Hessian matrix $\operatorname*{Hess}(r)(y)$ is invertible. The index of a nondegenerate routing point $y$ is the number of positive eigenvalues of $\operatorname*{Hess}(r)(y)$. Non-zero eigenvectors of $\operatorname*{Hess}(r)(y)$ to positive eigenvalues are called unstable eigenvector directions.
@@
    The norms of $r$, $\nabla r$, and $\operatorname*{Hess}(r)$ are bounded on ${\mathbb{R}}^{n}\smallsetminus{\mathcal{H}}$.
@@ 2.2 Gradient flow and connectivity via routing functions
Input: A routing function $r$ on ${\mathbb{R}}^{n}$ with numerator $h$, and its routing points $Z(r)=\{y_{1},\ldots,y_{m}\}$.
Output: A partition $\{C_{1},\dots,C_{s}\}$ of the set of routing points $Z(r)$ of $r$ on ${\mathbb{R}}^{n}$ corresponding to the connected components of ${\mathbb{R}}^{n}\smallsetminus V(h)$.
1ex
(1) Initialize a graph. Construct a graph $G$ with vertex set $Z(r)$ and no edges.
(2) Add edges via unstable directions.
@@
 foreach unstable eigenvector direction $v$ for $\operatorname*{Hess}(r)(y_{j})$ do
      Compute the limit routing point from $y_{j}$ in the direction $v$ with respect to $r$, say $y_{w_{+}}$.
      Add the edge $\{y_{j},y_{w_{+}}\}$ to $G$.
      Compute the limit routing point from $y_{j}$ in the direction v-v with respect to $r$, say $y_{w_{-}}$.
      Add the edge $\{y_{j},y_{w_{-}}\}$ to $G$.
(3) Transitive closure. Let $C_{1},\dots,C_{s}$ be the connected components of $G$.
@@ 2.3 Homotopy continuation and monodromy solving
Let $F(a,b,z)$ be as in Example 1.1. We wish to intersect the curve $V(F)\subseteq{\mathbb{C}}^{3}$ with the plane V(b1)V(b-1) by solving $F(x_{1},1,x_{2})=0$. For this, we can consider the homotopy
$H(x,t)=(1-t)F(x_{1},1,x_{2})+\gamma\,t\begin{pmatrix}x_{1}^{2}-1\\[2.0pt] x_{2}-1\end{pmatrix}=0,$
@@
The start points $(1,1)$ and (1,1)(-1,1) are nonsingular solutions of $H(x,1)=0$. Both corresponding paths are trackable and converge respectively to (2,1)(2,-1) and (2,1)(-2,1) of $H(x,0)=0$. This shows that the intersection of the zero set of $F$ with the plane V(b1)V(b-1) is $\{(2,1,-1),(-2,1,1)\}$.
@@
Recall that in Example 2.4, we computed the intersection of the curve $V(F)$ with the plane V(b1)V(b-1) by solving $F(x_{1},1,x_{2})=0$. We can vary that plane in a monodromy loop: For a loop $\Gamma\colon[0,1]\to{\mathbb{C}}$ with $\Gamma(0)=\Gamma(1)$, we consider now the new homotopy $H(x,t)=F(x_{1},\Gamma(t),x_{2})=0$ with start point (2,1)(2,-1). If $\Gamma(t)=3/4+(1/4)e^{2\pi i(1-t)}$ the corresponding path is trackable and returns to (2,1)(2,-1). In contrast, for the loop $\Gamma(t)=1/4+(3/4)e^{2\pi i(1-t)}$ the path ends at (2,1)(-2,1). Hence, the monodromy action exchanges these two solutions. The reason why the second loop gives a new solution, while the first does not, is that the second loop encircles a so-called branch point.
@@ 2.4 Summary of the algorithm for known defining polynomial
(3) Compute regions. Run Algorithm 1.
@@ 3 The case of an unknown defining polynomial
where $\pi\colon{\mathbb{C}}^{k}\times{\mathbb{C}}^{n-k}\to{\mathbb{C}}^{k}$ is the projection $\pi(p,z)=p$, and the line denotes Zariski closure. In what follows, we will let $h$ be a polynomial of minimal total degree such that $\mathcal{H}=V(h)$ (this defines $h$ uniquely up to scaling).
The key observation is that in order to carry out Algorithm 2, we do not need to have an explicit expression for $h$. Instead, we just need the following:
@@
  • We need to be able to evaluate $\nabla h$ and $\operatorname*{Hess}(h)$ in order to do homotopy continuation in step (2) and gradient flow in step (3).
It turns out that both these pieces of information about $h$ are available through a pseudo-witness set of $\mathcal{H}$. We explain this concept and how it gives the degree in Section 3.1. We then proceed to derive algorithms for evaluating $\nabla r$ and $\operatorname*{Hess}(r)$ through the data of a pseudo-witness set in Section 3.2, which is the main theoretical contribution of the paper. Building on this, we give a strategy (Algorithm 4) for computing ${\mathbb{R}}^{k}\smallsetminus\mathcal{H}$ in Section 3.3.
@@
It is often useful to consider the complement of the union of $\mathcal{H}$} with another hypersurface $V(g)$ where $g$ is a known polynomial, e.g., when imposing positivity conditions [17, 16]. This situation can be handled by simply adding $\log(g)$ to $\log(r)$ above and adjusting $e$ to ensure that $2e>\deg(h)+\deg(g)$.
@@ 3.1 Witness and pseudo-witness sets
It is important to note that the defining equations for $X$ are not used in Definition 3.2; instead, $X$ is represented using $F$. This can potentially introduce multiplicities, which is why we have the second item in the definition. The second item can be rephrased as follows: A witness set for an irreducible component $X$ of $V(F)$ is reduced when the ideal in the primary decomposition of $\langle F\rangle\vcentcolon=\langle f_{1},\ldots,f_{r}\rangle$ corresponding to $X$ is radical. In a reduced witness set, the points in $X\cap\mathcal{M}$ are nonsingular solutions to a square polynomial system in such a way that homotopy continuation can be used to trace the witness set to another witness set $X\cap\mathcal{M}^{\prime}$ for a different slice $\mathcal{M}^{\prime}$. This way, reduced witness sets enable several numerical operations in homotopy continuation, including sampling points on a variety, testing membership, and computing intersections [4, 31, 33, 49, 50, 51]. If a witness set $(F,\mathcal{M},{\mathcal{W}})$ is not reduced, one can utilize deflation to replace $F$ with a new witness system $F^{\prime}$ so that $(F^{\prime},\mathcal{M},{\mathcal{W}})$ is reduced [32].
@@
We give an example of a non-reduced witness-set. Consider the case $n=2$ and the polynomial $F(x_{1},x_{2})=(x_{1}^{2}+x_{2}^{2}-1)^{2}$. The zero set $X=V(F)$ is the circle. It has dimension $d=1$. The intersection of $V(F)$ with a general line consists of $2$ points. Take, for instance, the line $\mathcal{M}=V(x_{1})$. Then, $\mathcal{W}=X\cap\mathcal{M}=\{(0,1),(0,-1)\}$. However, $JF(x_{1},x_{2})=4(x_{1}^{2}+x_{2}^{2}-1)\,(x_{1},x_{2})$, so that $JF(0,1)=(0,0)$ and JF(0,1)=(0,0)JF(0,-1)=(0,0) both have rank 0, and not nd=21=1n-d=2-1=1. The witness set $(F,\mathcal{M},\mathcal{W})$ is therefore not reduced. This is caused by the fact that $x_{1}^{2}+x_{2}^{2}-1$ appears with multiplicity 2 in $F$.
@@
Projecting the coordinates of points in $\mathcal{W}\subset{\mathbb{C}}^{n}$ to the first $k$ coordinates yields $\pi(\mathcal{W})={\mathcal{H}}\cap\mathcal{L}$. Thus, we can compute the intersection of ${\mathcal{H}}\subset{\mathbb{C}}^{k}$ with a general line $\mathcal{L}$ by computing intersection of $X$ with $\mathcal{L}\times\mathcal{L}^{\prime}$. This yields the notion of pseudo-witness set, which we summarize in the following definition.
@@
    Let $X\subset{\mathbb{C}}^{n}$ be an irreducible variety of dimension $d$, such that ${\mathcal{H}}=\overline{\pi(X)}$ is a hypersurface. A pseudo-witness set for $\mathcal{H}$} is a quadruple $(F,\pi,\mathcal{L}\times\mathcal{L}^{\prime},\mathcal{W})$, where $F$ is a system such that $X$ is an irreducible component of $V(F)$, $\mathcal{L}$ is a general line and $\mathcal{L}^{\prime}$ a general linear space of dimension nd1n-d-1, and $\mathcal{W}=X\cap(\mathcal{L}\times\mathcal{L}^{\prime})$.
@@
Since $\pi(\mathcal{W})={\mathcal{H}}\cap\mathcal{L}$, the intersection of the discriminant with $\mathcal{L}$ is given by taking the first two coordinates of these points. The pseudo-witness set $(F,\pi,\mathcal{L}\times{\mathbb{C}},\mathcal{W})$ is reduced.
@@
Figure 3: The left picture illustrates a pseudo-witness set for the purple projection $\mathcal{H}$} of the red curve $V(F)$. While in Figure 2, the green linear space $\mathcal{M}$ was general, here it must be a product space: $\mathcal{M}=\mathcal{L}\times\mathbb{C}$, where $\mathcal{L}$ is a general line; see Example 3.6. The pseudo-witness set is given by the intersection $\mathcal{W}=V(F)\cap(\mathcal{L}\times{\mathbb{C}})$, and projection yields $\mathcal{H}$$\cap\mathcal{L}$. The right picture shows the situation within the yellow plane.